OpenStudy (anonymous):

find the value in simplest form:

OpenStudy (anonymous):

\[\sqrt{24} \over 4\sqrt{6}\]

OpenStudy (anonymous):

√24? @parthkohli?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {\sqrt{24} \times \sqrt{6} \over 24 } }\) \(\Large \color{Black}{\Rightarrow \sqrt{2 \times 2 \times 2 \times 3} \times \sqrt{6} \over 24 }\) \(\Large \color{Black}{\Rightarrow {2\sqrt{6} \times \sqrt{6} \over 24} }\) \(\Large \color{Black}{\Rightarrow {12 \over 24} }\) \(\Large \color{Black}{\Rightarrow 0.5 }\)

OpenStudy (anonymous):

lol I was so close o.o

OpenStudy (anonymous):

its not an option parth :l

OpenStudy (anonymous):

would it be 1?

OpenStudy (anonymous):

o wait nevermind theres 1/2 that will work right?

Parth (parthkohli):

1/2 = 9.5 lol

Parth (parthkohli):

0.5*

OpenStudy (anonymous):

9.5 o-o

Parth (parthkohli):

lol it's a typo